pH ammonium

1.     I have a beaker with 250.0 mL of distilled water.  I dissolve 10.0 g of ammonium chloride (NH­4Cl) and 10.0 g of methylammonium Chloride (CH3NH3Cl).  What is the pH of the resulting solution?
Answer #1

pH=pKa + log(b/a)

b=10/53.491

a=10/111.969

pH=5.11+ log(0.18/0.089)=5.43

source: instructor answer key
answered by: Tommy Cain